# ThinkEdit: Interpretable Weight Editing to Mitigate Overly Short Thinking in Reasoning Models

### v2

@@ Abstract
Recent studies have shown that Large Language Models (LLMs) augmented with chain-of-thought (CoT) reasoning demonstrate impressive problem-solving abilities. However, in this work, we identify a recurring issue where these models occasionally generate overly short reasoning, leading to degraded performance on even simple mathematical problems. Specifically, we investigate how reasoning length is embedded in the hidden representations of reasoning models and its impact on accuracy. Our analysis reveals that reasoning length is governed by a linear direction in the representation space, allowing us to induce overly short reasoning by steering the model along this direction. Building on this insight, we introduce ThinkEdit, a simple yet effective weight-editing approach to mitigate the issue of overly short reasoning. We first identify a small subset of attention heads (approximately 2%) that predominantly drive short reasoning behavior. We then edit the output projection weights of these heads to suppress the short reasoning direction. With changes to only 0.1% of the model’s parameters, ThinkEdit effectively reduces overly short reasoning and yields notable accuracy gains for short reasoning outputs (+5.44%), along with an overall improvement across multiple math benchmarks (+2.43%). Our findings provide new mechanistic insights into how reasoning length is controlled within LLMs and highlight the potential of fine-grained model interventions to improve reasoning quality. Our code is available at https://github.com/Trustworthy-ML-Lab/ThinkEdit.
@@ 1 Introduction
Recently, Reinforcement Learning (RL) has been applied to enhance Large Language Models (LLMs), equipping them with strong chain-of-thought (CoT) reasoning abilities [5]. These models, often referred to as reasoning models, first generate an intermediate reasoning process—a "thinking step"—where they reason step-by-step and then self-correct before producing a final response. As a result, they achieve remarkable improvement on mathematical reasoning tasks and demonstrate a strong ability to generate detailed CoT reasoning [9, 5, 15].
However, despite these improvements, reasoning models still exhibit a non-negligible gap from perfect accuracy on relatively simple benchmarks such as GSM8K [2]. As shown in Section 2, we found that Deepseek-distilled reasoning models occasionally generate overly short reasoning chains, which correlate with lower accuracy (about 20% drop on MATH-level5 benchmark [7]). This issue appears consistently across models of different sizes, suggesting that reasoning length plays a crucial role in problem-solving effectiveness. Yet, the mechanisms governing reasoning length within the model’s internal representation remain underexplored, despite being crucial for understanding reasoning models.
To bridge this gap, in this work, we first investigate how reasoning length is encoded within the hidden representations of reasoning models. By performing a novel analysis of the residual stream, we extract a reasoning length direction—a latent linear representation in the residual stream that enables direct control over reasoning length as shown in Figure 1 (left). Our analysis reveals that overly short, abstract, or high-level reasoning significantly degrades model performance, and this characteristic is primarily embedded in the middle layers of the model. Furthermore, we identify a small subset (approximately 2%) of attention heads in the middle layers that disproportionately contribute to short reasoning. Building on this insight, we propose ThinkEdit, a simple and effective weight-editing technique to remove the short-reasoning component from these attention heads’ output projection layers, as shown in Figure  1 (right). Our findings demonstrate that disabling these components leads to a non-trivial improvement in accuracy when the model generates short reasoning while also enhancing overall performance. Our contributions are summarized as follows:
@@
  • We discover a small set of "short reasoning" attention heads that strongly contribute to the generation of brief reasoning chains and propose ThinkEdit. By editing just 2% of the output projection weights of these heads (0.1% of the model’s total parameters), ThinkEdit effectively mitigates short reasoning, leading to improved performance both when short reasoning occurs (+5.44%) and in overall model accuracy (+2.43%).
@@
2 Unexpectedly Low Accuracy in Short Reasoning Cases
We begin our study by highlighting a consistent issue observed in Deepseek-distilled reasoning models across a variety of sizes: significantly lower accuracy when the reasoning length is short. This pattern holds across datasets such as GSM8K [2] and MATH-Level5 [7]. Figure 2 illustrates this trend, with the x-axis indicating a cutoff threshold on reasoning length. For example, a threshold of 2000 denotes that we calculate the average accuracy over all responses whose reasoning length is at most 2000 tokens. The y-axis shows the corresponding cumulative accuracy. The details of the experimental setup are provided in Section 4.4.
Contrary to intuition, one might expect shorter reasoning to correspond to easier questions, as such problems should require fewer steps to solve. This expectation is partially supported by the trend in Figure 2 (right), where accuracy tends to decrease as reasoning length exceeds 2000. However, the region with reasoning length below 2000 exhibits a different pattern: models consistently underperform on these short-reasoning cases, with accuracy dropping significantly below the overall average. This suggests that, rather than efficiently solving simple problems with brief reasoning, models often fail when producing overly short chains of thought.
Figure 2: Cumulative accuracy as a function of the reasoning length threshold. The x-axis represents the cutoff threshold on reasoning length, and the y-axis shows the corresponding cumulative accuracy. Models consistently exhibit lower accuracy for overly shorter reasoning (e.g. length <1000).
3 Understanding How Hidden Representations Affect Reasoning Length
@@ 3.1 Background of Transformer Structure and Notations
$r^{\text{attn}}_{\ell}=r^{\text{mlp}}_{\ell-1}+\textrm{Attn}(\textrm{LayerNorm}(r^{\text{mlp}}_{\ell-1})),\quad r^{\text{mlp}}_{\ell}=r^{\text{attn}}_{\ell}+\textrm{MLP}(\textrm{LayerNorm}(r^{\text{attn}}_{\ell}))$
@@
The models analyzed in this study, derived from LLaMA3.1 [4] and Qwen2.5 [25], adopt a pre-normalization Transformer architecture, where layer normalization is applied before both the self-attention and MLP modules. Our focus is on the hidden representations $r^{\text{attn}}_{\ell}$ and $r^{\text{mlp}}_{\ell}$ as illustrated in Figure 1 (left), which capture the model’s state after the self-attention and MLP transformations, respectively.
@@ 3.2 Extracting Reasoning Length Directions
To investigate how reasoning length is encoded in a model’s hidden representation, we begin by collecting the model’s responses to 1,000 problems from the GSM8K [2] training set. In each response, the chain-of-thought (CoT) is enclosed between special tags <think> and </think>. We measure the length of each CoT by counting only the tokens within these tags. Based on this length, we construct two datasets $\mathcal{D}_{\text{long}}$ and $\mathcal{D}_{\text{short}}$, where $\mathcal{D}_{\text{long}}$ consists of responses whose CoT exceeds 1000 tokens and $\mathcal{D}_{\text{short}}$ includes those under 100 tokens. Each entry in these datasets contains: (1) the problem statement, (2) the extracted CoT, enclosed by <think> and </think> tags, and (3) the step-by-step calculation process leading to the final answer.
@@
$\overline{r}^{\text{attn}}_{\ell,\text{long}}=\frac{1}{|\mathcal{D}_{\text{long}}|}\sum_{i\in\mathcal{D}_{\text{long}}}\frac{1}{|\mathcal{T}_{i}|}\sum_{t\in\mathcal{T}_{i}}r^{\text{attn}}_{\ell}(i,t),\quad\overline{r}^{\text{attn}}_{\ell,\text{short}}=\frac{1}{|\mathcal{D}_{\text{short}}|}\sum_{i\in\mathcal{D}_{\text{short}}}\frac{1}{|\mathcal{T}_{i}|}\sum_{t\in\mathcal{T}_{i}}r^{\text{attn}}_{\ell}(i,t),$
$\overline{r}^{\text{mlp}}_{\ell,\text{long}}=\frac{1}{|\mathcal{D}_{\text{long}}|}\sum_{i\in\mathcal{D}_{\text{long}}}\frac{1}{|\mathcal{T}_{i}|}\sum_{t\in\mathcal{T}_{i}}r^{\text{mlp}}_{\ell}(i,t),\quad\overline{r}^{\text{mlp}}_{\ell,\text{short}}=\frac{1}{|\mathcal{D}_{\text{short}}|}\sum_{i\in\mathcal{D}_{\text{short}}}\frac{1}{|\mathcal{T}_{i}|}\sum_{t\in\mathcal{T}_{i}}r^{\text{mlp}}_{\ell}(i,t).$
Finally, we define the reasoning-length direction at layer $\ell$ as the vector difference between the “long” and “short” embeddings:
$v^{\text{attn}}_{\ell}\;=\;\overline{r}^{\text{attn}}_{\ell,\text{long}}\;-\;\overline{r}^{\text{attn}}_{\ell,\text{short}},\quad\quad v^{\text{mlp}}_{\ell}\;=\;\overline{r}^{\text{mlp}}_{\ell,\text{long}}\;-\;\overline{r}^{\text{mlp}}_{\ell,\text{short}}.$ (1)
@@ 3.3 Effects of Reasoning-Length Direction
$r^{\text{attn}}_{\ell}\;\leftarrow\;r^{\text{attn}}_{\ell}\;+\;\alpha\,v^{\text{attn}}_{\ell},\quad\quad r^{\text{mlp}}_{\ell}\;\leftarrow\;r^{\text{mlp}}_{\ell}\;+\;\alpha\,v^{\text{mlp}}_{\ell}.$ (2)
@@
We evaluate the effect of these reasoning-length directions—derived from 1,000 GSM8K training examples—using two test sets:
  • GSM8K (200 problems) [2]: A relatively simpler benchmark, consisting of the first 200 problems from the GSM8K test set.
  • MATH-Level5 (140 problems) [7]: A more challenging benchmark, comprising 140 problems selected from the MATH test set. Specifically, we extract 20 level-5 (highest difficulty) examples from each of 7 categories.
We set a maximum reasoning length of 8,192 tokens for GSM8K and 16,384 tokens for MATH-Level5. Upon reaching this limit, the model is prompted to finalize its answer immediately. We experiment on three publicly available reasoning models on Huggingface of different sizes: deepseek-distill-qwen-1.5B, deepseek-distill-llama3-8B, and deepseek-distill-qwen-14B.
@@
Figure 3 illustrates our global steering results on GSM8K. The top row applies the attention-based direction $v_{\ell}^{\text{attn}}$, while the bottom row applies the MLP-based direction $v_{\ell}^{\text{mlp}}$. We vary $\alpha$ from $-0.08$ (shorter reasoning) to $+0.08$ (longer reasoning). Across all models, increasing $\alpha$ extends the length of CoT (Figure 3, right), indicating that $v_{\ell}^{\text{attn}}$ and $v_{\ell}^{\text{mlp}}$ indeed encode reasoning-length attributes. In terms of accuracy, the larger 8B and 14B models improve when steered toward longer reasoning—particularly deepseek-distill-llama3-8B (orange line), which benefits most from positive steering with 10% accuracy improvement. In contrast, the smaller deepseek-distill-qwen-1.5B (blue line) model experiences a 10% drop in accuracy. Figure 4 presents the results for the more challenging MATH-Level5 dataset. Similar to GSM8K, our extracted directions effectively control reasoning length as expected, with negative $\alpha$ consistently leading to shorter CoT and positive $\alpha$ extending them. In terms of accuracy, shorter reasoning also consistently degrades performance. However, unlike GSM8K, there is no clear trend indicating that longer reasoning reliably enhances accuracy; while moderate positive $\alpha$ might provide some benefits, excessively long reasoning often negatively impacts performance.
Figure 3: Global steering on GSM8K. Each row shows a different steering vector (top: $v_{\ell}^{\text{attn}}$, bottom: $v_{\ell}^{\text{mlp}}$). Positive $\alpha$ extends reasoning length across all models and improves accuracy in the 8B and 14B models, whereas negative $\alpha$ consistently shortens reasoning and lowers accuracy across all models.
Figure 4: Global steering on MATH-Level5. Negative $\alpha$ shortens reasoning and lowers accuracy.
@@
We conduct a layerwise experiment to determine which layers are most influenced by the reasoning-length direction, as detailed in Appendix A.1. Our findings indicate that the middle layers exhibit the most significant fluctuations, suggesting they play a crucial role in encoding representations that control reasoning length.
@@
Recent work [15] proposed an interesting approach to enforce budget constraints by stopping the CoT or appending "Wait" to prolong it. However, stopping the CoT prematurely may cause incomplete reasoning and appending "Wait" may risk misalignment with the model’s natural CoT. Alternatively, steering representations may allow for a more coherent way to modulate reasoning length (see Section A.2) – by directly manipulating the model’s internal representations, one can more effectively balance the computational cost and performance.
@@
Based on these findings, we hypothesize that certain critical components within the middle layers may contribute to short reasoning. In the next section, we adopt a mechanistic interpretability approach to pinpoint these components and perform weight editing to mitigate their effects.
@@ 4.1 Overview of Attention-Head Structure
  • Q, K, and V Projections. Given a hidden-state ${r}\in\mathbb{R}^{T\times d}$ (for $T$ tokens), each head $h$ computes:
    ${Q}^{h}={r}{W}_{q}^{h}\;\in\;\mathbb{R}^{T\times d_{h}},\quad{K}^{h}={r}{W}_{k}^{h}\;\in\;\mathbb{R}^{T\times d_{h}},\quad{V}^{h}={r}{W}_{v}^{h}\;\in\;\mathbb{R}^{T\times d_{h}}.$
    Each head $h$ has its own learnable projection matrices ${W}_{q}^{h},{W}_{k}^{h},{W}_{v}^{h}\in\mathbb{R}^{d\times d_{h}}$, which transform the hidden representation ${r}$ into query, key, and value vectors.
@@
    ${A}^{h}=\mathrm{softmax}\!\Bigl{(}\tfrac{{Q}^{h}({K}^{h})^{\top}}{\sqrt{d_{h}}}\Bigr{)}\,{V}^{h}\;\;\in\;\mathbb{R}^{T\times d_{h}}.$
@@
    $C^{h}\coloneq{A}^{h}{W}_{o}^{h}\;\in\;\mathbb{R}^{T\times d}.$ (3)

### v3

@@ Abstract
Recent studies have shown that Large Language Models (LLMs) augmented with chain-of-thought (CoT) reasoning demonstrate impressive problem-solving abilities. However, in this work, we identify a recurring issue where these models occasionally generate overly short reasoning, leading to degraded performance on even simple mathematical problems. Specifically, we investigate how reasoning length is embedded in the hidden representations of reasoning models and its impact on accuracy. Our analysis reveals that reasoning length is governed by a linear direction in the representation space, allowing us to induce overly short reasoning by steering the model along this direction. Building on this insight, we introduce ThinkEdit, a simple yet effective weight-editing approach to mitigate the issue of overly short reasoning. We first identify a small subset of attention heads (approximately 4%) that predominantly drive short reasoning behavior. We then edit the output projection weights of these heads to remove the short reasoning direction. With changes to only 0.2% of the model’s parameters, ThinkEdit effectively reduces overly short reasoning and yields notable accuracy gains for short reasoning outputs (+6.39%), along with an overall improvement across multiple math benchmarks (+3.34%). Our findings provide new mechanistic insights into how reasoning length is controlled within LLMs and highlight the potential of fine-grained model interventions to improve reasoning quality. Our code is available at: https://github.com/Trustworthy-ML-Lab/ThinkEdit
ThinkEdit: Interpretable Weight Editing to
Mitigate Overly Short Thinking in Reasoning Models
Chung-En Sun UCSD CSE cesun@ucsd.edu                        Ge Yan UCSD CSE geyan@ucsd.edu                        Tsui-Wei Weng UCSD HDSI lweng@ucsd.edu
@@ 1 Introduction
Recently, Reinforcement Learning (RL) has been applied to enhance Large Language Models (LLMs), equipping them with strong chain-of-thought (CoT) reasoning abilities Guo et al. (2025). These models, often referred to as reasoning models, first generate an intermediate reasoning process—a "thinking step"—where they reason step-by-step and then self-correct before producing a final response. As a result, they achieve remarkable improvement on mathematical reasoning tasks and demonstrate a strong ability to generate detailed CoT reasoning Jaech et al. (2024); Guo et al. (2025); Muennighoff et al. (2025).
However, despite these improvements, reasoning models still exhibit a non-negligible gap from perfect accuracy on relatively simple benchmarks such as GSM8K Cobbe et al. (2021). As shown in Section 2, we found that Deepseek-distilled reasoning models occasionally generate overly short reasoning chains, which correlate with lower accuracy (about 20% drop on MATH-level5 benchmark Hendrycks et al. (2021b)). This issue appears consistently across models of different sizes, suggesting that reasoning length plays a crucial role in problem-solving effectiveness. Yet, the mechanisms governing reasoning length within the model’s internal representation remain underexplored, despite being crucial for understanding reasoning models.
Figure 1: Cumulative accuracy as a function of the reasoning length threshold. The x-axis represents the cutoff threshold on reasoning length, and the y-axis shows the average accuracy of all responses with reasoning length below that threshold. Models consistently exhibit lower accuracy for overly short reasoning (e.g. length <1000).
To bridge this gap, in this work, we first investigate how reasoning length is encoded within the hidden representations of reasoning models. By performing a novel analysis of the residual stream, we extract a reasoning length direction—a latent linear representation in the residual stream that enables direct control over reasoning length as shown in Figure 2 (left). Our analysis reveals that overly short, abstract, or high-level reasoning significantly degrades model performance, and this characteristic is primarily embedded in the middle layers of the model. Furthermore, we identify a small subset (approximately 4%) of attention heads in the middle layers that disproportionately contribute to short reasoning. Building on this insight, we propose ThinkEdit, a simple and effective weight-editing technique to remove the short-reasoning component from these attention heads’ output projection layers, as shown in Figure  2 (right). Our findings demonstrate that disabling these components leads to a non-trivial improvement in accuracy when the model generates short reasoning while also enhancing overall performance. Our contributions are summarized as follows:
@@
  • We discover a small set of "short reasoning" heads that strongly contribute to the generation of brief reasoning chains and propose ThinkEdit. By editing the output projection weights of just 4% heads (0.2% of the model’s total parameters), ThinkEdit effectively mitigates short reasoning, leading to improved performance both when short reasoning occurs (+6.39%) and in overall accuracy (+3.34%).
2 Unexpectedly Low Accuracy in Short Reasoning Cases
We begin our study by highlighting a consistent issue observed in Deepseek-distilled reasoning models across a variety of sizes: significantly lower accuracy when the reasoning length is short. This pattern holds across datasets such as GSM8K Cobbe et al. (2021) and MATH-Level5 Hendrycks et al. (2021b). Figure 1 illustrates this trend, with the x-axis indicating a cutoff threshold on reasoning length. For example, a threshold of 2000 denotes that we calculate the average accuracy over all responses whose reasoning length is at most 2000 tokens. The y-axis shows the corresponding cumulative accuracy. The details of the experimental setup are provided in Section 4.4.
Contrary to intuition, one might expect shorter reasoning to correspond to easier questions, as such problems should require fewer steps to solve. This expectation is partially supported by the trend in Figure 1 (right), where accuracy tends to decrease as reasoning length exceeds 2000. However, the region with reasoning length below 2000 (highlighted in red boxes) exhibits a different pattern: models consistently underperform on these short-reasoning cases, with accuracy dropping significantly below the overall average. This suggests that, rather than efficiently solving simple problems with brief reasoning, models often fail when producing overly short chains of thought.
@@ 2 Unexpectedly Low Accuracy in Short Reasoning Cases
3 Understanding How Representations Affect Reasoning Length
@@ 3.1 Background of Transformer Structure and Notations
$r^{\text{attn}}_{\ell}$ $=r^{\text{mlp}}_{\ell-1}+\textrm{Attn}(\textrm{LayerNorm}(r^{\text{mlp}}_{\ell-1}))$
$r^{\text{mlp}}_{\ell}$ $=r^{\text{attn}}_{\ell}+\textrm{MLP}(\textrm{LayerNorm}(r^{\text{attn}}_{\ell}))$
@@
Our focus is on the hidden representations $r^{\text{attn}}_{\ell}$ and $r^{\text{mlp}}_{\ell}$ as illustrated in Figure 2 (left), which capture the model’s state after the self-attention and MLP transformations, respectively.
@@ 3.2 Extracting Reasoning Length Directions
To investigate how reasoning length is encoded in a model’s hidden representation, we begin by collecting the model’s responses to 2,000 problems from the GSM8K Cobbe et al. (2021) training set. In each response, the chain-of-thought (CoT) is enclosed between special tags <think> and </think>. We measure the length of each CoT by counting only the tokens within these tags. We then construct two datasets $\mathcal{D}_{\text{long}}$ and $\mathcal{D}_{\text{short}}$, where $\mathcal{D}_{\text{long}}$ consists of responses whose CoT exceeds 1000 tokens and $\mathcal{D}_{\text{short}}$ includes those under 100 tokens. Each entry in these datasets contains: (1) the problem statement, (2) the extracted CoT, enclosed by <think> and </think> tags, and (3) the step-by-step calculation process leading to the final answer.
@@
$\overline{r}^{x}_{\ell,y}=\frac{1}{|\mathcal{D}_{y}|}\sum_{i\in\mathcal{D}_{y}}\frac{1}{|\mathcal{T}_{i}|}\sum_{t\in\mathcal{T}_{i}}r^{x}_{\ell}(i,t),\vspace{-5pt}$
where $x\in\{\text{attn},\text{mlp}\}$ denotes the representation type and $y\in\{\text{long},\text{short}\}$ indicates the reasoning-length group. Finally, we define the reasoning-length direction at layer $\ell$ as the vector difference between the “long” and “short” embeddings:
$v^{\text{attn}}_{\ell}=\overline{r}^{\text{attn}}_{\ell,\text{long}}-\overline{r}^{\text{attn}}_{\ell,\text{short}},\;\;v^{\text{mlp}}_{\ell}=\overline{r}^{\text{mlp}}_{\ell,\text{long}}-\overline{r}^{\text{mlp}}_{\ell,\text{short}}.$ (1)
@@
Figure 3: Global steering results. Top: On GSM8K, positive $\alpha$ extends reasoning length and improves accuracy in the 8B and 14B models, while negative $\alpha$ shortens reasoning and lowers accuracy. Bottom: On MATH-Level5, negative $\alpha$ similarly shortens reasoning and reduces accuracy.
@@ 3.3 Effects of Reasoning-Length Direction
$r^{\text{attn}}_{\ell}\leftarrow r^{\text{attn}}_{\ell}+\alpha\,v^{\text{attn}}_{\ell},\quad r^{\text{mlp}}_{\ell}\leftarrow r^{\text{mlp}}_{\ell}+\alpha\,v^{\text{mlp}}_{\ell}.\vspace{-5pt}$ (2)
@@
We evaluate the effect of reasoning-length directions using two test sets:
  • GSM8K (200 problems) Cobbe et al. (2021): A simpler benchmark, consisting of the first 200 problems from the GSM8K test set.
  • MATH-Level5 (140 problems) Hendrycks et al. (2021b): A more challenging benchmark, comprising 140 problems selected from the MATH test set. Specifically, we extract 20 level-5 examples from each of 7 categories.
We set a maximum reasoning length of 8,192 tokens for GSM8K and 16,384 tokens for MATH-Level5. Upon reaching this limit, the model is prompted to finalize its answer immediately. We experiment on three reasoning models of varying sizes: deepseek-distill-{qwen-1.5B, llama3-8B, qwen-14B}.
@@
Figure 3 (Top) shows the effect of applying the attention-based direction $v_{\ell}^{\text{attn}}$ on GSM8K. We vary $\alpha$ from $-0.08$ (shorter reasoning) to $+0.08$ (longer reasoning). Across all models, increasing $\alpha$ extends the length of CoT (Figure 3, top right), indicating that $v_{\ell}^{\text{attn}}$ indeed encode reasoning-length attributes. In terms of accuracy, the larger 8B and 14B models improve when steered toward longer reasoning—particularly deepseek-distill-llama3-8B (orange line), which benefits most from positive steering with 10% accuracy improvement. In contrast, the smaller deepseek-distill-qwen-1.5B (blue line) model experiences a 10% drop in accuracy. Figure 3 (Bottom) presents the results for the more challenging MATH-Level5 dataset. Similar to GSM8K, our extracted directions effectively control reasoning length as expected, with negative $\alpha$ consistently leading to shorter CoT and positive $\alpha$ extending them. In terms of accuracy, shorter reasoning also consistently degrades performance. However, unlike GSM8K, there is no clear trend indicating that longer reasoning reliably enhances accuracy; while moderate positive $\alpha$ might provide some benefits, excessively long reasoning often negatively impacts performance. We also present results using the MLP-based direction $v_{\ell}^{\text{mlp}}$ in Appendix A.1, which exhibit similar trends.
@@
We perform a layerwise experiment to identify which layers produce reasoning-length directions with the strongest impact. As shown in Appendix A.2, middle layers are most effective, suggesting their key role in encoding reasoning-length representations.
@@
Recent work Muennighoff et al. (2025) proposed an interesting approach to enforce budget constraints by stopping the CoT or appending "Wait" to prolong it. However, stopping the CoT prematurely may cause incomplete reasoning and appending "Wait" may risk misalignment with the model’s natural CoT. Alternatively, steering representations may allow for a more coherent way to modulate reasoning length (see Appendix A.6) – by directly manipulating the model’s internal representations, one can more effectively balance the computational cost and performance.
@@
Based on these findings, we hypothesize that certain critical components within the middle layers may contribute to short reasoning. In the next section, we pinpoint these components and perform weight editing to mitigate their effects.
@@ 4.1 Overview of Attention-Head Structure
  • Q, K, and V Projections. Given a hidden-state ${r}\in\mathbb{R}^{T\times d}$ (for $T$ tokens), each head $h$ computes: $Q^{h}=rW_{q}^{h},\;K^{h}=rW_{k}^{h},\;V^{h}=rW_{v}^{h},$ where $Q^{h},K^{h},V^{h}\!\in\!\mathbb{R}^{T\times d_{h}}$. Each head $h$ has its own learnable projection matrices ${W}_{q}^{h},{W}_{k}^{h},{W}_{v}^{h}\!\in\!\mathbb{R}^{d\times d_{h}}$, which transform the hidden representation ${r}$ into query, key, and value vectors.
@@
    ${A}^{h}=\mathrm{softmax}\!\Bigl{(}\tfrac{{Q}^{h}({K}^{h})^{\top}}{\sqrt{d_{h}}}\Bigr{)}\,{V}^{h}\;\;\in\;\mathbb{R}^{T\times d_{h}}.\vspace{-5pt}$
@@
    $C^{h}\coloneq{A}^{h}{W}_{o}^{h}\;\in\;\mathbb{R}^{T\times d}.\vspace{-5pt}$ (3)